OpenStudy (anonymous):

The function f(t) = 20 sin (pi over 5t) + 12 models the temperature of a periodic chemical reaction where t represents time in hours. What are the maximum and minimum temperatures of the reaction, and how long does the entire cycle take? Maximum: 20°; minimum: 8°; period: 12 hours Maximum: 32°; minimum: −8°; period: 10 hours Maximum: 20°; minimum: 12°; period: pi over 5 hours Maximum: 32°; minimum: 8°; period: pi over 5 hours @mathmale

OpenStudy (anonymous):

is it c?

OpenStudy (mathmale):

I will always ask you to explain / show how you arrived at any particular answer. What clues have you found that seem to indicate that the answer is c?

OpenStudy (anonymous):

i just looked at what was give. @mathmale

OpenStudy (mathmale):

Important: Your f(t) = 20 sin (pi over 5t) + 12 should be written as f(t) = 20 sin (pi over 5)t + 12, I believe. Agree or disagree?

OpenStudy (anonymous):

ik the period would be pi over 5.

OpenStudy (mathmale):

So, what is your argument that the answer is c?

OpenStudy (mathmale):

Look at your f(t) = 20 sin (pi over 5)t + 12. Compare this to f(t) = a sin bt + d The value of b here is pi over 5. The period is 2pi/b. Knowing b, find the period now, please.

OpenStudy (anonymous):

that the a is the max and that is 20 and the 12 is the min and the min is d

OpenStudy (anonymous):

is the max and min right?

OpenStudy (mathmale):

'a' is the AMPLITUDE of your sine function. What is the maximum value that the given sine function will attain? The minimum? Explain how you got these figures. 'd' is not the minimum. 'd' represents your VERTICAL OFFSET.

OpenStudy (mathmale):

If any of these terms are not familiar to you, either look them up or ask me specific questions.

OpenStudy (anonymous):

i am so confused!! how do i find the max and the min in this problem?

OpenStudy (mathmale):

Suppose that t=0. What would the value of your sine function be? Just substitute 0 for t in f(t) = 20 sin (pi over 5)t + 12.

OpenStudy (anonymous):

12

OpenStudy (mathmale):

Right. 12 = 'd' = your vertical offset. Good. Now add 20 to 12. What do you get? Now subtract 20 from 12. What do you get? What do these two results represent?

OpenStudy (anonymous):

20+12=32 and 20-12=8

OpenStudy (anonymous):

ohhh wow so the answer is a

OpenStudy (mathmale):

Add 20 to 12: you get 20+12=32. Yes. Subtract 20 from 12. Instead, you have subtracted 12 from 20. No. Try again: Subtract 20 from 12.

OpenStudy (anonymous):

my b i meant d

OpenStudy (anonymous):

ohh 12-20=-8

OpenStudy (mathmale):

OK. In summary, adding the amplitude, 'a,' or 20, to the vertical offset, 'd,' or 12, results in 32. What does 32 represent? subtracting the amplitude, 'a,' or 20, from the vertical offset, 'd,' or 12, results in -8. what does -32 represent?

OpenStudy (anonymous):

32 is the maximum

OpenStudy (mathmale):

Yes, and ... ?

OpenStudy (anonymous):

-8 is the minimum

OpenStudy (mathmale):

Nice work. Please review our conversation (above) and determine what 'b' is in f(t) = 20 sin (pi over 5)t + 12 .

OpenStudy (anonymous):

so the period is 10

OpenStudy (mathmale):

Yes. Don't mind me... but please explain how you got 10.

OpenStudy (anonymous):

5 times 2 and the pies cancel

OpenStudy (mathmale):

Good. So, which of the four possible answers is the correct one? Is it c?

OpenStudy (anonymous):

no its not c, it is b

OpenStudy (mathmale):

Right. Now you see why I rejected your "Is it c?" question. Wanted to be absolutely sure that you knew how to find a, b and d and what to do with them to find the min, the max and the period. Nice work.

OpenStudy (anonymous):

thank you so much! :-) can you help me with a couple more?

OpenStudy (mathmale):

Perhaps one more. Have to get off the Internet soon. Always please post a problem with which you want help before asking for help. Thanks.

OpenStudy (mathmale):

You're welcome; happy to help you!

OpenStudy (anonymous):

ok and thanks again.

OpenStudy (anonymous):

(cos Θ − cos Θ)2 + (cos Θ + cos Θ)2